Given f (x) = x +1 and g(x) = 2x − 3, determine (f*g) (11/2)?
Answer by Theo(13342)

Given f(x) = (x+1) and g(x) = (2x-3), find f*g(11/2).

since (f*g)(x) = f(x) * g(x), then:

(f*g)(x) = (x+1) * (2x-3).

if you are looking for (f*g)(11/2), the you just replace x with 11/2 and you get:

(f*g)(11/2) = (11/2+1) * (2(11/2)-3) which becomes:

(f*g)(11/2) = (13/2) * (16/2) = 52

you could also have solved it using x and then evaluated it using x = 11/2.

(f*g)(x) = (x+1)*(2x-3) = 2x^2 -3x + 2x - 6 = 2x^2 - x - 3.

when x = 11/2, this becomes 2*(11/2)^2 - (11/2) - 3 which is equal to 52.